# Arithmetic aspects of discrete periodic Toda flows

### v2

@@ 2.6 Mumford’s algebraic description of $\mathrm{Jac(\mathcal{C})}$
2.6.1 Mumford’s representation
Recall that any reduced divisor $D^{\prime}$ not supported at infinity can be represented by two polynomials $P(x),Q(x)\in R[x]$, such that $P$ is monic, $\mathrm{deg}(P)\leq g$ and
$P\ |\ Q^{2}+hQ+f.$ (2.74)
@@ 2.6.1 Mumford’s representation
$(\alpha,Q(\alpha))\in\mathcal{C}.$ (2.75)
@@
$D^{\prime}=[P,Q\text{ mod }P].$ (2.76)
2.6.2 All of $\mathrm{Jac(\mathcal{C})}$
Let us introduce the divisor $D_{\infty}=\lceil\frac{g}{2}\rceil\infty^{+}+\lfloor\frac{g}{2}\rfloor\infty^{-}$. Following [5], any element of $\mathrm{Jac(\mathcal{C})}$ can be uniquely written as difference $D-D_{\infty},$ with
$D=D^{\prime}+n\infty^{+}+(g-n-\mathrm{deg}(D^{\prime}))\infty^{-},$ (2.77)
where $D^{\prime}$ is a Mumford representation and $0\leq n\leq g-\mathrm{deg}(D^{\prime})$ is a natural number. This is the so-called balanced (Mumford) representation. With the notation from above, we obtain the following description
$\mathrm{Jac(\mathcal{C})}=\{([P,Q],n)\}.$ (2.78)
When the balancing parameter $n$ is understood, we simply write $[P,Q]=([P,Q],n)\in\mathrm{Jac(\mathcal{C})}$.
2.6.3 Examples
@@
$[0]=([1,0],\lceil\tfrac{g}{2}\rceil)\in\mathrm{Jac(\mathcal{C})}.$ (2.79)
@@
$\mathfrak{pa}=[\infty^{+}-\infty^{-}]=([1,0],\lceil\tfrac{g}{2}\rceil+1)\in\mathrm{Jac(\mathcal{C})},$ (2.80)
$-\mathfrak{pa}=[\infty^{-}-\infty^{+}]=([1,0],\lceil\tfrac{g}{2}\rceil-1)\in\mathrm{Jac(\mathcal{C})}$ (2.81)
will appear prominently in the following. The element $\mathfrak{pa}$ corresponds to a fundamental solution of the famous Pell-Abel equation for $\mathcal{C}$, as explained in [6]. Our framework gives a natural proof of the fact that $\mathfrak{pa}\in\mathrm{Jac(\mathcal{C})}[n]$, see Theorem 2.1.
2.7 Gauß composition and addition law in $\mathrm{Jac(\mathcal{C})}$
@@
$([P_{1},Q_{1}],n_{1})\boxplus([P_{2},Q_{2}],n_{2})=([P,Q],n)\in\mathrm{Jac(\mathcal{C})}$ (2.82)
in our set-up. The main ingredient is Cantor’s [^6] Originally, Cantor worked over imaginary hyperelliptic curves where all subtleties at infinity disappear, in our set-up of real hyperelliptic curves some care is initially needed. adaption [1] of Gauß famous composition law for quadratic forms (over $\mathbb{Z}$) adapted to hyperelliptic curves (see also [17]), which consists of a composition step and reduction step for the Mumford representations. We extract the part of the general algorithm for real hyperelliptic curves from [5] necessary for our set-up as follows.
@@ I. Composition
$s=\mathrm{gcd}(P_{1},P_{2},Q_{1}+Q_{2}+h)=f_{1}P_{1}+f_{2}P_{2}+f_{3}(Q_{1}+Q_{2}+h).$ (2.83)
Then the composition step $[\tilde{P},\tilde{Q}]$ is defined by
$[\tilde{P},\tilde{Q}]=[P_{1}P_{2}/s^{2},(f_{1}P_{1}Q_{2}+f_{2}P_{2}Q_{1}+f_{3}(Q_{1}Q_{2}-f))/s\text{ mod }\tilde{P}].$ (2.84)
@@ II. Reduction
The reduction step $[P,Q]=[P_{1},Q_{1}]\boxplus[P_{2},Q_{2}]$ is defined by
$[P,Q]=[(\tilde{Q}^{2}+h\tilde{Q}+f)/\tilde{P}\text{ made monic},-(\tilde{Q}+h)\text{ mod }P].$ (2.85)
@@ III. Composition at infinity
We will only look at sums of the form $([u,v],0)\boxplus\mathfrak{a}$, where $\mathfrak{a}\in\{\mathfrak{pa},\mathfrak{T}\}$ (see (2.119) for the definition of the latter). In those two cases, the set-up works such that the result is of the form $([\overline{u},\overline{v}],0)$ where $[\overline{u},\overline{v}]$ is given by the composition of the Mumford representations.
2.7.1 Example of inverse element
@@
$-([P,Q],n)=\left\{\begin{array}[]{c}([P,-(h+Q)\text{ mod }P],g-n-\mathrm{deg}(P)+1)\text{, if $n>0$}\\ \ ([P,-(h+Q)\text{ mod }P],0)\boxminus\mathfrak{pa}\text{, if $n=0$}\end{array}.\right.$ (2.87)
2.8 The eigenvector map $\Psi$ and its inverse
@@
$\phi=\begin{bmatrix}\phi_{1}(x,z)\\ \vdots\\ \phi_{n}(x,z)\end{bmatrix},$ (2.88)
@@
$\mathcal{L}\,\phi=\begin{bmatrix}0\\ \vdots\\ 0\\ (-1)^{n}|\mathcal{L}|\end{bmatrix},$ (2.90)
i.e., on the spectral curve $\mathcal{C}$ the vector $\phi$ is an eigenvector of $L_{t}(z)$ with eigenvalue $-x$. In particular we immediately see
$\phi_{n}(x,z)=0\iff u(x)=0,$ (2.91)
$\phi_{1}(x,z)=0\iff v(x)=z,$ (2.92)
$\phi_{n-1}(x,z)=0\iff w(x)=f/z.$ (2.93)
@@
$\mathcal{R}_{\mathcal{C}}=\{(\mathbb{I},\mathbb{V})\in R^{2n}\ |\ \mathcal{C}_{(\mathbb{I},\mathbb{V})}=\mathcal{C}\}.$ (2.94)
@@
$\Psi_{\mathcal{C}}:\mathcal{R}_{\mathcal{C}}\longrightarrow\mathrm{Jac(\mathcal{C})}$ (2.95)
@@
$(\mathbb{I},\mathbb{V})\longmapsto([u,v],0).$ (2.96)
@@
The degree of $u$ is equal to $g=n-1$ and the degree of $v$ is equal to $g-1$. The condition
$u\,|\,v^{2}+vh+f$ (2.97)
@@
$I_{n}=\frac{v_{x}+I_{o}}{u_{x}}\ \ \ \text{ and }\ \ \ V_{n-1}=\frac{w_{x}+V_{o}}{u_{x}}.$ (2.99)
@@
$\Psi^{-1}_{\mathcal{C}}:\mathrm{Im}(\Psi_{\mathcal{C}})\longrightarrow\mathcal{R}_{\mathcal{C}}.$ (2.104)
@@
$I_{k}=\tfrac{\sigma^{k}(v)_{x}+I_{o}}{\sigma^{k}(u)_{x}}$ (2.105)
@@
$V_{k}=\tfrac{\mathrm{l.c.}(\sigma^{k}(v))}{I_{k}}.$ (2.106)
@@ 2.8.1 Further properties of $\Psi$
$[\sigma(u),\sigma(v))]=[u,v]\boxplus\mathfrak{pa}\in\mathrm{Jac(\mathcal{C})}.$ (2.107)
@@
$\mathfrak{pa}\in\mathrm{Jac}(\mathcal{C})[n].$ (2.108)
@@
$[\sigma(u),\sigma(v)]=[(v^{2}+vh+f)/(I_{n}V_{n}u),(I_{1}+V_{n}+x)\sigma(u)-(h+v)].$ (2.109)
For the first equality, we recall that
$x\sigma(u)=h_{I_{1},V_{n}}=(I_{n}+V_{n-1}+x)u_{I_{1},V_{n}}-w_{I_{1},V_{n}}.$ (2.110)
@@
$\sigma(u)=(I_{n}+V_{n-1}+x)|{}^{1}\mathcal{L}_{1}|-I_{n-1}V_{n-1}|{}^{1}\mathcal{L}_{2}|-.$ (2.113)
@@
$(v^{2}+vh+f)/(I_{n}V_{n}u)=(I_{n}+V_{n-1}+x)|{}^{1}\mathcal{L}_{1}|-I_{n-1}V_{n-1}|{}^{1}\mathcal{L}_{2}|.$ (2.114)
For the second equality we use $\sigma(h)=h$ and $\sigma(w)=v$. ∎
@@
$-[u,v]=\left\{\begin{array}[]{c}[u,w]\text{ ,if $g$ even}\\ \ [u,w]\boxminus\mathfrak{pa}\text{, if $g$ odd}\end{array}.\right.$ (2.115)
@@
$[u,w]\boxplus\mathfrak{pa}=[\sigma^{-1}(u),\sigma^{-1}(w)].$ (2.116)
@@
$[u,w]\boxplus\mathfrak{pa}=\boxminus[u,v]\boxplus\mathfrak{pa}=\boxminus([u,v]\boxminus\mathfrak{pa})=\boxminus(\sigma^{-1}[u,v])=\sigma^{-1}[u,w].$ (2.117)
@@
$[u^{\dagger},v^{\dagger}]={[u,w]}\boxminus\mathfrak{pa}.$ (2.118)
@@
2.9 Linearization of the discrete periodic Toda flow
@@
$\mathfrak{T}=([x,-I_{o}],\lceil\tfrac{g}{2}\rceil-1)\in\mathrm{Jac}(\mathcal{C}).$ (2.119)
@@
$[\underline{u},\underline{v}]=[u,v]\boxplus\mathfrak{T}\in\mathrm{Jac(\mathcal{C})},$ (2.120)
@@
${R^{2n}}$${\mathrm{Jac}(\mathcal{C})}$${R^{2n}}$${\mathrm{Jac}(\mathcal{C}).}$$\scriptstyle{\Psi}$$\scriptstyle{\mathcal{T}}$$\scriptstyle{\cdot\boxplus\mathfrak{T}}$$\scriptstyle{\Psi}$ (2.121)
@@
$[\underline{u},\underline{v}]=[(\tilde{v}^{2}+\tilde{v}h+f)/(-I_{n}xu),(x+V_{n})\underline{u}-(h+\tilde{v})].$ (2.123)
@@
$\tilde{v}^{2}+\tilde{v}h+f=-I_{n}(V_{n-1}+x)u^{2}+I_{n}uw+(V_{n-1}+x)uv-I_{n-1}I_{n}V_{n-1}V_{n}u|{}^{1}\mathcal{L}_{2}|.$ (2.125)
Dividing by $-I_{n}u$ we obtain
$(\tilde{v}^{2}+\tilde{v}h+f)/(-I_{n}u)$ (2.126)
=\ = $(V_{n-1}+x)u-w-(V_{n-1}+x)v/I_{n}+I_{n-1}V_{n-1}V_{n}|{}^{1}\mathcal{L}_{2}|$ (2.127)
=\ = $(V_{n-1}+x)u_{V_{n}}-w_{V_{n}}$ (2.128)
=\ = $xu_{V_{n}}+V_{n-1}u_{I_{n-1},V_{n}}.$ (2.129)
@@
Thus showing the first equality. The second equality follows again from Lemma 2.6 and $h=\underline{h}$. ∎
@@
$-\mathfrak{T}=\left\{\begin{array}[]{c}([x,-V_{o}],\lfloor\tfrac{g}{2}\rfloor)\text{, if $g$ even}\\ \ ([x,-V_{o}],\lfloor\tfrac{g}{2}\rfloor+1)\text{, if $g$ odd}\end{array}\right.\in\mathrm{Jac}(\mathcal{C}).$ (2.131)
2.10 Discrete periodic Toda flow via the cyclic shift map
@@
$x\underline{u}$ $=xu+V_{n-1}u_{I_{n-1}}+v-V_{n}\sigma(u)$ (2.132)
$=h+2v-I_{n}u-V_{n}\sigma(u),$ (2.133)
$\underline{v}$ $=V_{n}\underline{u}+v-V_{n}\sigma(u).$ (2.134)
@@
$x\underline{u}$ $=(\tilde{v}^{2}+\tilde{v}h+f)/(-I_{n}u)$ (2.135)
$=(\tilde{v}((V_{n-1}+x)u-w)+f)/(-I_{n}u)$ (2.136)
$=((-I_{n}u+v)(V_{n-1}+x)u-(-I_{n}u+v)w+f)/(-I_{n}u)$ (2.137)
$=(V_{n-1}+x)u-w+((V_{n-1}+x)uv-vw+f)/(-I_{n}u).$ (2.138)
@@
$x\underline{u}$ $=h-\tilde{v}-\tilde{v}^{\dagger},$ (2.142)
$\sigma(u)$ $=\underline{u}-(\underline{v}-v)/V_{n}=\underline{u}_{I_{1}}+v/V_{n}$ (2.143)
$=((I_{n}+x)v+V_{n-1}v_{I_{n-1}})/(I_{n}V_{n}).$ (2.144)
@@
2.11 Cantor’s division polynomials
We want to explain the following theorem due to Cantor
Theorem 2.3 ([2] Theorem 8.7).
$r\cdot[X-x,y]=[\Delta_{r}(\tfrac{a-X}{4b^{2}},E_{r}(\tfrac{a-X}{4b^{2}})]\in\mathrm{Jac(\mathcal{D})}.$ (2.145)
2.11.1 Discrete open Toda flow
Set $E(z)=F(a-z)$ and consider the formal power series
$S(z)=\sum_{i\geq 0}s_{i}z^{i}=(-1)^{g+1}\sqrt{E(z)}.$ (2.146)
Define the Hankel determinants
$h_{m,n}=(-1)^{\binom{n}{2}}|(s_{m+i-j})_{i,j=1}^{n}|.$ (2.147)
$h_{m-1,n}h_{m+1,n}=h_{m,n-1}h_{m,n+1}+h_{m,n}^{2}.$ (2.148)
Remark 2.11.
The relation (2.148) is equivalent to the open discrete Toda equation in the $\tau$-function formalism, as explained in [7, 4].
2.11.2 Cantor’s division polynomials
[…]
$f_{r-1}A^{i}_{r+1}(z)=f_{r}A^{i}_{r}(z)-zf_{r+1}A^{i}_{r-1}(z)$ (2.150)
with initial conditions
$(A^{1}_{g-1},A^{1}_{g})=(-z^{g-1},-z^{g})\text{ and }(A^{2}_{g-1},A^{2}_{g})=(0,0).$ (2.151)
$D_{r}(z)=-(A^{1}_{r}(z)^{2}-A^{2}_{r}(z)^{2}E(z))/z^{r}$ (2.152)
and
$\tilde{E}(z)=\tfrac{zf_{r-1}f_{r+1}}{2f_{r}^{2}}(\tfrac{D_{r+1}(z)}{f_{r+1}^{2}}-\tfrac{D_{r-1}(z)}{f_{r-1}^{2}})\text{ mod }D_{r}(z).$ (2.153)
$\Delta_{r}(z)=(2b)^{2\nu_{r}}D_{r}(4b^{2}z)\text{ and }E_{r}(z)=\tilde{E}_{r}(4b^{2}z),$ (2.154)
where $\nu_{r}=\binom{r}{2}-\binom{g}{2}.$
2.11.3 Back to our set-up
$t\cdot\mathfrak{T}\in\mathrm{Jac(\mathcal{C})}.$ (2.155)
$\mathfrak{T}_{\alpha}=[x_{\alpha}-\tfrac{1}{\alpha},\tfrac{V_{o}-I_{o}}{\beta(-\alpha)^{g+1}}]\in\mathrm{Jac}(\mathcal{C}_{\alpha})$ (2.156)
to which we can apply Cantor’s division polynomials (for $t>g$).
Remark 2.12.
As far as the author knows, there is surprisingly no direct (algebraic) approach to division polynomials for real hyperelliptic curves in the literature. Later we will see that $p$-adic methods might pave the way to such a theory.
3 How to recover the periodic box-ball flow

### v3

@@ 2.6. Mumford’s algebraic description of $\mathrm{Jac(\mathcal{C})}$
2.6.1. Mumford’s representation
Note any reduced divisor $D^{\prime}$ not supported at infinity can be represented by two polynomials $P(x),Q(x)\in R[x]$, such that $P$ is monic, $\mathrm{deg}(P)\leq g$ and
$P\ |\ Q^{2}+hQ+f.$
@@ 2.6.1. Mumford’s representation
$(\alpha,Q(\alpha))\in\mathcal{C}.$
@@
$D^{\prime}=[P,Q\text{ mod }P].$
2.6.2. All of $\mathrm{Jac(\mathcal{C})}$
We introduce the divisor $D_{\infty}=\lceil\frac{g}{2}\rceil\infty^{+}+\lfloor\frac{g}{2}\rfloor\infty^{-}$. Following [5], any element of $\mathrm{Jac(\mathcal{C})}$ can be uniquely written as difference $D-D_{\infty},$ with
$D=D^{\prime}+d\infty^{+}+(g-d-\mathrm{deg}(D^{\prime}))\infty^{-},$
where $D^{\prime}$ is a Mumford representation and $0\leq d\leq g-\mathrm{deg}(D^{\prime})$ is a natural number. This is the so-called balanced (Mumford) representation. Using the above notation, we obtain the following description
$\mathrm{Jac(\mathcal{C})}=\{([P,Q],d)\}.$
When the balancing parameter $d$ is understood, we simply write $P,Q]=([P,Q],d)\in\mathrm{Jac(\mathcal{C})}$.
2.6.3. Examples
@@
$0]=([1,0],\lceil\tfrac{g}{2}\rceil)\in\mathrm{Jac(\mathcal{C})}.$
@@
$\mathfrak{pa}=[\infty^{+}-\infty^{-}]=([1,0],\lceil\tfrac{g}{2}\rceil+1)\in\mathrm{Jac(\mathcal{C})},$
$\mathfrak{pa}=[\infty^{-}-\infty^{+}]=([1,0],\lceil\tfrac{g}{2}\rceil-1)\in\mathrm{Jac(\mathcal{C})}$
will appear prominently in the following. The element $\mathfrak{pa}$ corresponds to a fundamental solution of the famous Pell–Abel equation for $\mathcal{C}$, as explained in [6]. Our framework gives a natural proof of the fact that $\mathfrak{pa}\in\mathrm{Jac(\mathcal{C})}[n]$, see Theorem 2.1.
2.7. Gauß composition and addition law in $\mathrm{Jac(\mathcal{C})}$
@@
$([P_{1},Q_{1}],d_{1})\boxplus([P_{2},Q_{2}],d_{2})=([P,Q],d)\in\mathrm{Jac(\mathcal{C})}$
in our set-up. The main ingredient is Cantor’s [^5] Originally, Cantor worked over imaginary hyperelliptic curves where all subtleties at infinity disappear, in our set-up of real hyperelliptic curves some care is initially needed. adaptation [1] of Gauß composition law for quadratic forms (over $\mathbb{Z}$) to hyperelliptic curves (see also [17]), which consists of a composition step and reduction step for the Mumford representations. We extract the part of the general algorithm for real hyperelliptic curves from [5] necessary for our set-up as follows.
@@ I. Composition
$s=\mathrm{gcd}(P_{1},P_{2},Q_{1}+Q_{2}+h)=f_{1}P_{1}+f_{2}P_{2}+f_{3}(Q_{1}+Q_{2}+h).$
Then the composition step $\tilde{P},\tilde{Q}$] is defined by
[…]
@@ II. Reduction
The reduction step $P,Q]=[P_{1},Q_{1}]\boxplus[P_{2},Q_{2}]$ is defined by
[…]
@@ III. Composition at infinity
We will only look at sums of the form $([u,v],0)\boxplus\mathfrak{a}$, where $\mathfrak{a}\in\{\mathfrak{pa},\mathfrak{T}\}$ (see (2.7) for the definition of the latter). In those two cases, the set-up works such that the result is of the form $([\overline{u},\overline{v}],0)$ where $\overline{u},\overline{v}]$ is given by the composition of the Mumford representations.
2.7.1. Example of inverse element
@@
[…]
2.8. The eigenvector map $\Psi$ and its inverse
@@
$\phi=\begin{bmatrix}\phi_{1}(x,z)\\ \vdots\\ \phi_{n}(x,z)\end{bmatrix}$,
@@
$\mathcal{L}\,\phi=\begin{bmatrix}0\\ \vdots\\ 0\\ (-1)^{n}|\mathcal{L}|\end{bmatrix}$,
i.e., on the spectral curve $\mathcal{C}$ the vector $\phi$ is an eigenvector of $L_{t}$ with eigenvalue x-x. In particular we immediately see
(2.5) $\phi_{n}(x,z)=0\iff u(x)=0,$
(2.6) $\phi_{1}(x,z)=0\iff v(x)=z,$
$\phi_{n-1}(x,z)=0\iff w(x)=f/z.$
@@
$\mathcal{R}_{\mathcal{C}}=\{(\mathbb{I},\mathbb{V})\in R^{2n}\ |\ \mathcal{C}_{(\mathbb{I},\mathbb{V})}=\mathcal{C}\}.$
@@
$\Psi_{\mathcal{C}}:\mathcal{R}_{\mathcal{C}}\longrightarrow\mathrm{Jac(\mathcal{C})}$
@@
$(\mathbb{I},\mathbb{V})\longmapsto([u,v],0).$
@@
The degree of $u$ is equal to g=n1g=n-1 and the degree of $v$ is equal to g1g-1. The condition
$u\,|\,v^{2}+vh+f$
@@
$I_{n}=\frac{v_{x}+I_{\Pi}}{u_{x}}\ \ \ \text{ and }\ \ \ V_{n-1}=\frac{w_{x}+V_{\Pi}}{u_{x}}.$
@@
$\Psi^{-1}_{\mathcal{C}}:\mathrm{Im}(\Psi_{\mathcal{C}})\longrightarrow\mathcal{R}_{\mathcal{C}}.$
@@
$I_{k}=\tfrac{\sigma^{k}(v)_{x}+I_{\Pi}}{\sigma^{k}(u)_{x}}$
@@
$V_{k}=\tfrac{\mathrm{l.c.}(\sigma^{k}(v))}{I_{k}}.$
@@ 2.8.1. Further properties of $\Psi$
$u^{\sigma},v^{\sigma}]=[u,v]\boxplus\mathfrak{pa}\in\mathrm{Jac(\mathcal{C})}.$
@@
$\mathfrak{pa}\in\mathrm{Jac}(\mathcal{C})[n].$
@@
$u^{\sigma},v^{\sigma}]=[(v^{2}+vh+f)/(I_{n}V_{n}u),(I_{1}+V_{n}+x)u^{\sigma}-(h+v)].$
For the first equality, we recall
$xu^{\sigma}=h_{I_{1},V_{n}}=(I_{n}+V_{n-1}+x)u_{I_{1},V_{n}}-w_{I_{1},V_{n}}.$
@@
$u^{\sigma}=(I_{n}+V_{n-1}+x)|{}^{1}\mathcal{L}_{1}|-I_{n-1}V_{n-1}|{}^{1}\mathcal{L}_{2}|.$
@@
$(v^{2}+vh+f)/(I_{n}V_{n}u)=(I_{n}+V_{n-1}+x)|{}^{1}\mathcal{L}_{1}|-I_{n-1}V_{n-1}|{}^{1}\mathcal{L}_{2}|.$
For the second equality we use $h^{\sigma}=h$ and $w^{\sigma}=v$. ∎
@@
$u,v]=\left\{\begin{array}[]{c}[u,w]\text{ ,if $g$ even}\\ \ [u,w]\boxminus\mathfrak{pa}\text{, if $g$ odd}\end{array}.\right$.
@@
$u,w]\boxplus\mathfrak{pa}=[\sigma^{-1}(u),\sigma^{-1}(w)].$
@@
$u,w]\boxplus\mathfrak{pa}=\boxminus[u,v]\boxplus\mathfrak{pa}=\boxminus([u,v]\boxminus\mathfrak{pa})=\boxminus(\sigma^{-1}[u,v])=\sigma^{-1}[u,w].$
@@
$u^{\dagger},v^{\dagger}]={[u,w]}\boxminus\mathfrak{pa}.$
@@
2.9. Linearization of the discrete periodic Toda flow
@@
(2.7) $\mathfrak{T}=([x,-I_{\Pi}],\lceil\tfrac{g}{2}\rceil-1)\in\mathrm{Jac}(\mathcal{C}).$
@@
$\overline{u},\overline{v}]=[u,v]\boxplus\mathfrak{T}\in\mathrm{Jac(\mathcal{C})},$
@@
$R^{2n}$$\mathrm{Jac}(\mathcal{C})$$R^{2n}$$\mathrm{Jac}(\mathcal{C}).$$\scriptstyle{\Psi}$$\scriptstyle{\mathcal{T}}$$\scriptstyle{\cdot\boxplus\mathfrak{T}}$$\scriptstyle{\Psi}$
@@
[…]
@@
$\tilde{v}^{2}+\tilde{v}h+f=-I_{n}(V_{n-1}+x)u^{2}+I_{n}uw+(V_{n-1}+x)uv-I_{n-1}I_{n}V_{n-1}V_{n}u|{}^{1}\mathcal{L}_{2}|.$
Dividing by $I_{n}u$ we obtain
$(\tilde{v}^{2}+\tilde{v}h+f)/(-I_{n}u)$
$=\$$(V_{n-1}+x)u-w-(V_{n-1}+x)v/I_{n}+I_{n-1}V_{n-1}V_{n}|{}^{1}\mathcal{L}_{2}|$
$=\$$(V_{n-1}+x)u_{V_{n}}-w_{V_{n}}$
[…]
@@
Thus showing the first equality. The second equality follows again from Lemma 2.6 and $h=\overline{h}$. ∎
@@
[…]
2.10. Discrete periodic Toda flow via the cyclic shift map
@@
$x\overline{u}$ $=xu+V_{n-1}u_{I_{n-1}}+v-V_{n}u^{\sigma}$
$=h+2v-I_{n}u-V_{n}u^{\sigma},$
$\overline{v}$ $=V_{n}\overline{u}+v-V_{n}u^{\sigma}.$
@@
$x\overline{u}$ $=(\tilde{v}^{2}+\tilde{v}h+f)/(-I_{n}u)$
$=(\tilde{v}((V_{n-1}+x)u-w)+f)/(-I_{n}u)$
$=((-I_{n}u+v)(V_{n-1}+x)u-(-I_{n}u+v)w+f)/(-I_{n}u)$
$=(V_{n-1}+x)u-w+((V_{n-1}+x)uv-vw+f)/(-I_{n}u).$
@@
$x\overline{u}$ $=h-\tilde{v}-\tilde{v}^{\dagger}$,
$u^{\sigma}$ $=\overline{u}-(\overline{v}-v)/V_{n}=\overline{u}_{I_{1}}+v/V_{n}$
$=((I_{n}+x)v+V_{n-1}v_{I_{n-1}})/(I_{n}V_{n}).$
@@
3. How to recover the periodic box-ball flow